# Bohao Wu   Qingyun Wang   Yue Guo

### v2

@@ B.3 Model Analysis in Terms of Jargon Domain
Conversely, in the ’Philosophy’ domain, NAb model faces challenges in both precision and recall. It exhibits a higher false positive rate, misclassifying terms such as ’Structural constraints’, ’Analytic philosophers’, and ’Argument Facets’. This indicates a tendency to over-identify common philosophical terms as highly specialized jargon. Furthermore, NAb model also demonstrates lower recall in the Philosophy domain, failing to detect several jargon terms, including ‘Computational argumentation’, ‘Corpus with 320 arguments’, ‘Nonmonotonic inference methods’, ‘Super-knotty rope’, ‘Super-knot’, and ‘Dialectical quality’. These terms represent complex philosophical concepts that the NAb model struggles to recognize as domain-specific jargon.

### v3

@@ B.3 Model Analysis in Terms of Jargon Domain
Conversely, in the ‘Philosophy’ domain, NAb model faces challenges in both precision and recall. It exhibits a higher false positive rate, misclassifying terms such as ‘Structural constraints’, ‘Analytic philosophers’, and ‘Argument Facets’. This indicates a tendency to over-identify common philosophical terms as highly specialized jargon. Furthermore, NAb model also demonstrates lower recall in the Philosophy domain, failing to detect several jargon terms, including ‘Computational argumentation’, ‘Corpus with 320 arguments’, ‘Nonmonotonic inference methods’, ‘Super-knotty rope’, ‘Super-knot’, and ‘Dialectical quality’. These terms represent complex philosophical concepts that the NAb model struggles to recognize as domain-specific jargon.